If income is $100, income tax is $10, consumption is $60, government tax revenue is $10 and government expenditure is $ 8, what is the domestic saving?

a. $30

b. $32

c. $40

d. $25

e. none of the above
Answer #1

Answer

Option b

$32

======

Domestic saving =Y-C-G

=100-60-8

=$32

======

or

Disposable income =income -tax=100-10=90

pricate Saving =income -consumption =90-60=30

Public saving =T-G=10-8=2

Sum of the savings =30+2=32

the domestic saving is $32

Option b